Calculate the mole fraction of a 20% aqueous hydrochloric acid solution

1
Expert's answer
2022-04-17T11:04:30-0400

20% HCl solution means

20 g of HCl in 100 g solution

m(H2O) = 100 - 20 = 80 g

n(HCl) = m(HCl) / Mr(HCl) = 20 g / 36.5 g/mol = 0.548 mol

n(H2O) = m(H2O) / Mr(H2O) = 80 g / 18 g/mol = 4.444 mol

n(solution) = 0.548 mol + 4.444 mol = 4.992 mol

"\chi"(HCl) = n(HCl) / n(solution) = 0.548 mol / 4.992 mol = 0.11